# Towards Better Comprehension of Breaking Changes in the NPM Ecosystem

### v1

@@ 8. Conclusion and Future Work
In this study, for better comprehension of breaking changes in the NPM ecosystem and enhancing BC-related tools, we conduct an empirical study to bridge the knowledge gap from three aspects, with our carefully constructed breaking change dataset (1,519 BCs in total) sampled from a large set of popular NPM projects. We find that 97% of BCs detected by regression testing can be covered by developers’ documentation, which proves that extracting BCs from documentation is reasonable. We then summarize the JavaScript and TypeScript specific syntactic BCs and how developers make behavioral BCs, yielding many findings. Besides, we also investigate the reasons behind BCs in JavaScript and conclude with a taxonomy, which extend the previous works on motivations behind breaking changes. Based on our empirical findings, we provide actionable implications for future research, e.g., applying automatic renaming and naming techniques in JavaScript projects, and detecting code with similar functionalities, etc.
In the future, we want to collect additional sources such as online discussions, and employ alternative approaches like dynamic analysis to gain a deeper understanding of breaking changes in the NPM ecosystem. We plan to investigate how to automatically identify the breaking changes in NPM projects. We consider combining static analysis and dynamic analysis techniques to enhance the existing BC detection approaches. We also consider improving BC detection with the help of Large Language Models since they show great performance improvements in many software engineering related tasks and can utilize the semantics in source code.

### v2

@@ 8. Conclusion and Future Work
In this study, for better comprehension of breaking changes in the NPM ecosystem and enhancing breaking change-related tools, we conducted an empirical study to bridge the knowledge gap from three aspects, with our carefully constructed breaking change dataset (1,519 breaking changes in total) sampled from a large set of popular NPM projects. We found that 95.4% of the breaking changes detected by regression testing can be covered by developers’ documentation, which proves that extracting breaking changes from documentation is reasonable. We then summarized the breaking changes in the NPM projects that are specific to JavaScript and TypeScript, and how developers make behavioral breaking changes, which yield many findings. Besides, we also investigated the reasons behind breaking changes in JavaScript and conclude with a taxonomy, which extend the previous works on motivations behind breaking changes. Based on our empirical findings, we provided actionable implications for future research, e.g., applying automatic renaming and naming techniques in JavaScript projects, and detecting code with similar functionalities, etc.
In the future, we want to collect additional sources such as online discussions, and employ alternative approaches like dynamic analysis to gain a deeper understanding of breaking changes in the NPM ecosystem. We plan to investigate how to automatically identify the breaking changes in the NPM projects. We consider combining static analysis and dynamic analysis techniques to enhance the existing breaking change detection approaches. We also consider improving breaking change detection with the help of large language models since they show great performance improvements in many software engineering related tasks and can utilize the semantics in source code.
Acknowledgements.
This research/project is supported by the National Science Foundation of China (No.62372398, No.72342025, and U20A20173), the Fundamental Research Funds for the Central Universities (No. 226-2022-00064), and the National Research Foundation, under its Investigatorship Grant (NRF-NRFI08-2022-0002). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore.